OpenStudy (anonymous):

8z^2-64 when factored out does this equal 8(z-8)(z+8)

OpenStudy (anonymous):

8 ( z^2 - 8 )

OpenStudy (dumbcow):

when you factor out the 8 you need to divide the 64 by 8 as well

OpenStudy (amistre64):

its just a matter of taking the time to multiply it all together again; you can do it :)

OpenStudy (anonymous):

so when i multiply it all back out it gives me my 8z^2-64 ohhhh ok lol im just trying to get other peoples opinions o my answers and check my answers for an assighnment i have

OpenStudy (amistre64):

(2z sqrt(2) +8) (2z sqrt(2) -8) is what i get

OpenStudy (amistre64):

8.8.8 > 64

OpenStudy (anonymous):

we arent supposed to use the quadratic equation we dont learn that til junior year

OpenStudy (amistre64):

this factors to a difference of squares

OpenStudy (amistre64):

your original conclusion there is wrong

OpenStudy (anonymous):

but when i multiply it all back it gives me the original equation

OpenStudy (anonymous):

?

OpenStudy (anonymous):

what would it equal then cause its confusing me

OpenStudy (anonymous):

\[= 8 ( z^2 -8 )\]\[= 8 ( z-\sqrt{8})(z+\sqrt{8})\]

OpenStudy (anonymous):

thats it , noting else

OpenStudy (anonymous):

or you could write it as \[8 ( z-2\sqrt{2})(z+2\sqrt{2})\] as \[\sqrt{8} = \sqrt{4 \times 2} = \sqrt{4} \times \sqrt{2} = 2\sqrt{2}\]